Three-point charges of 0.07uC and 0.24uC are 35 cm apart in air. What is the potential energy of the system?






1
Expert's answer
2022-02-28T13:40:46-0500

The potential energy of system


"W=k\frac{q_1q_2}{r}\\
=9*10^9*\frac{0.07*10^{-6}*0.24*10^{-6}}{0.35}=4.32*10^{-4}\;\rm J"